OpenStudy (anonymous):

What is the 20th term of the following arithmetic sequence? 1/2, 1 , 3/2, 2... @jim_thompson5910

OpenStudy (jdoe0001):

wht's the "common difference"?

OpenStudy (anonymous):

a+19d

OpenStudy (anonymous):

idk fractions that well.

OpenStudy (anonymous):

The common difference is 1/2

OpenStudy (anonymous):

@bingtamekia okay, so what do i do next?

OpenStudy (anonymous):

apply the formular given to you = a+19d

OpenStudy (anonymous):

where a=1/2 and d=1/2

OpenStudy (anonymous):

so i guess you will have 10 as your answer

OpenStudy (anonymous):

anyone agree? i just wanna make sure

OpenStudy (anonymous):

Yeah you will get 10

OpenStudy (anonymous):

thank you all!

OpenStudy (anonymous):

np :D